For the following reaction:



2NaCl(s) + H2SO4(g) → Na2SO4(s) + 2HCl(g)


Compound    ΔH°f (kJ mol-1)    S° (J mol-1 K-1)
NaCl (s)    -411.12    72.11
H2SO4 (g)    -735.13    298.78
Na2SO4 (s)    -1387.1    149.6
HCl (g)    -92.31    186.90





Calculate ΔG°rx (in kJ) at 120.21 K for this reaction. Report your answer to two decimal places in standard notation (i.e. 123.45 kJ).

Assume ΔH°f and S° do not vary as a function of temperature.
Answer #1

Step 1:

Given:

Hof(NaCl(s)) = -411.12 KJ/mol

Hof(H2SO4(g)) = -735.13 KJ/mol

Hof(Na2SO4(s)) = -1387.1 KJ/mol

Hof(HCl(g)) = -92.31 KJ/mol

Balanced chemical equation is:

2 NaCl(s) + H2SO4(g) ---> Na2SO4(s) + 2 HCl(g)

ΔHo rxn = 1*Hof(Na2SO4(s)) + 2*Hof(HCl(g)) - 2*Hof( NaCl(s)) - 1*Hof(H2SO4(g))

ΔHo rxn = 1*(-1387.1) + 2*(-92.31) - 2*(-411.12) - 1*(-735.13)

ΔHo rxn = -14.35 KJ

Step 2:

Given:

Sof(NaCl(s)) = 72.11 J/mol.K

Sof(H2SO4(g)) = 298.78 J/mol.K

Sof(Na2SO4(s)) = 149.6 J/mol.K

Sof(HCl(g)) = 186.9 J/mol.K

Balanced chemical equation is:

2 NaCl(s) + H2SO4(g) ---> Na2SO4(s) + 2 HCl(g)

ΔSo rxn = 1*Sof(Na2SO4(s)) + 2*Sof(HCl(g)) - 2*Sof( NaCl(s)) - 1*Sof(H2SO4(g))

ΔSo rxn = 1*(149.6) + 2*(186.9) - 2*(72.11) - 1*(298.78)

ΔSo rxn = 80.4 J/K

Step 3:

ΔHo = -14.35 KJ/mol

ΔSo = 80.4 J/mol.K

= 0.0804 KJ/mol.K

T = 120.21 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -14.35 - 120.21 * 0.0804

ΔGo = -24.01 KJ

Answer: -24.01 KJ